# TULIP: Transformer for Upsampling of LiDAR Point Clouds

### v2

@@ 3.2.1 Patch Partition
The patch partition layer creates an initial feature embedding from the input range image. This initial feature representation significantly influences the network’s performance. Specifically, the selection of this layer determines how the inherent characteristics, patterns, and relevant information of the LiDAR range image data are encoded into a format that the network can effectively learn from. Most transformer-based RGB image super-resolution approaches utilize a relatively large patch size to tokenize the input [26, 9, 31]. These larger patches enable the construction of a global spatial context, facilitating an understanding of the overall structure and high-level features of the RGB image. These networks employ square-shaped patches and implement upsampling in both vertical and horizontal directions. In contrast, LiDAR upsampling is primarily aiming to enhance the vertical resolution of the input, given that range images possess properties that are vastly different from RGB images. Furthermore, attention across more pixels in range image is less useful than in RGB image, since there is almost no geometrical relation for the spatial contexts that are far from each other in the 2D range image space. Motivated by this, we propose two adjustments to effectively process range images within a Swin-Unet:
@@
Row-Based Patch Tokenization: Our model builds row patches with a dimension of $1\times 4$ for range image tokenization. The new patch geometry is designed to retain full vertical information while compressing horizontally. This aligns with our objective of conserving and extending vertical details in the range image. Besides that, row patches excel at capturing boundary discontinuities between distinct objects and the background within the scene. Moreover, the selection of the patch size strikes a balance between enhancing model performance and limiting model capacity. Smaller patches can capture more detailed local information, but they also increase the number of model parameters, which can slow down the model’s training process. Circular Padding (CD) in Horizontal Dimension: Instead of adding zero pixels padding, our model uses circular padding in the horizontal dimension. This avoids introducing artificial features and naturally matches the sensor model of a rotating 3D LiDAR. This preserves accurate neighbor relations along the edges, which would otherwise be compromised during the projection to the 2D range image. The Circular Padding (CD) also allows us to fit flexible input sizes without changing the patch geometry.
@@ 3.2.2 Non-Square window for local self-attention
Inside Swin Transformer, input feature maps are window-partitioned along width and height to compute the self-attention locally in each window. In addition to square-shaped RGB images, some recent work applies a square window for panoramic data as well [28, 52]. However, using square windows one range images is unfavorable. On one hand, small windows can focus more on attention within near areas but lead to increasing computational complexity and can struggle with the reconstruction of objects at a further distance in the scene. On the other hand, a larger square window can blend scene-related contexts at different scales, which can potentially degrade the network performance. In a prior study done by Eskandar et al. [17], they demonstrated the significant enhancement in range image super-resolution by separately upsampling the upper and lower portions of the data. In contrast to their approach of utilizing CNN-based shallower and deeper network branches for distinct feature extraction, our proposal involves the direct partitioning of the input using a rectangular window. This window geometry facilitates robust attention interaction among distant points through W-MSA and subtle cross-attention between objects at multiple scales via SW-MSA and hierarchical processing (refer to Fig. 4). In addition, range images, as typical panorama data, inherently contain abundant information along the horizontal direction. Therefore, directing attention more toward the image width rather than the height can aid the network in capturing a greater amount of geometrical information. At the same time, the computational complexity remains the same as using a large square window.
Figure 4: Our network uses a hierarchical structure for patch merging on the range image. We propose to use rectangle window to fit the range data and to handle depth discontinuity at different scales.
3.2.3 Additional Modifications and Postprocessing
Patch Unmerging (PU) and Pixel Shuffle (PS): Similar to the patch merging layer that is responsible for downsampling at the encoder stage, a mechanism is necessary within the decoder to upsample the feature maps. Uformer [45] achieves upsampling through a de-convolutional layer. As they tend to smooth out image sharpness [26], we design a reverse operation of patch merging. We re-arrange the channels of patches into their respective positions in a grid to upsample the resolution of input feature maps, which we denote as patch unmerging. Furthermore, we built the projection head based on the pixel shuffle operation [39]. Monte Carlo Dropout: We additionally use Monte Carlo Dropout (MC-Dropout) [38] to filter unreliable points. MC-Dropout executes several feed-forward passes during inference with different active dropouts, which yields a distribution of outputs. We refine the results by removing points with high output variance, as they often come down to noisy, invalid points.
@@ 4.1 Experimental Settings
Datasets: We conduct experiments on three different datasets, that include the two large-scale real-world datasets DurLAR [25] and KITTI [19] (Sec. 4.3.2), as well as on a dataset [24] that was generated using the CARLA simulator [15] (Sec. 4.3.1). For all datasets, we select test sequences that are recorded in different locations than the train set to avoid spatial overlap. We vertically downsample the high-resolution range images with a factor of four by skipping the respective lines to generate low-resolution input images. For ablation studies in Sec. 4.2, we select the KITTI dataset. Implementation Details: In all experiments, we use distributed processing with 4×4$\times$4 × GeForce RTX 2080 Ti. For training, we use a fixed batch size of eight per GPU for all datasets. We use the AdamW [30] optimizer with a weight decay factor of 0.05 and a base learning rate of 5e-4. Evaluation metrics: We evaluate the Mean Absolute Error (MAE) for all the pixels in the generated 2-D range images. Additionally, we assess the performance by considering the 3D points reconstructed by our neural networks. The Chamfer Distance (CD) measures the Euclidean distance between two point clouds. We also follow the approach in [24] to evaluate the volumetric occupancy similarity. To do so, the point clouds are voxelized using a voxel size of $0.1m$. A voxel is classified as occupied for each point cloud if it contains at least one point. We then calculate the Intersection-over-Unit (IoU) based on the occupancy.
@@ 4.2 Ablation Studies
Model 3 $1\times 4$ Rectangular $\times$ $\times$ 0.4228 0.4040 0.1218
Model 4 $1\times 4$ Rectangular $\checkmark$ $\times$ 0.4186 0.4094 0.1199
TULIP $1\times 4$ Rectangular $\checkmark$ $\checkmark$ 0.4183 0.4134 0.1198
TULIP-L $1\times 4$ Rectangular $\checkmark$ $\checkmark$ 0.3678 0.4329 0.0971
Table 1: Ablation study results for 4×4$\times$4 × upsampling on KITTI. Input Resolution: $16\times 1024$, Output Resolution: $64\times 1024$.
@@
Window geometry: To verify the effectiveness of modifying window geometry for range images, we conduct experiments with square and rectangular windows (Tab. 1 Model 2-3. Based on the quantitative results, we can observe the following: reconstruction quality produced by local attention within large square windows is inferior to the rectangular window with respect to both 2D and 3D evaluation metrics, while they result in the same computational complexity. The underlying reason for this performance difference lies in the perceptive field of the windows. A wider window is more likely to include discontinuities between objects at different scales. Accordingly, the network performs better at separating objects from each other, which can visually be validated by clearer surface boundaries in Fig. 6.
Additional Modifications and Postprocessing: We additionally apply Circular Padding (CP), Patch Unmerging (PU), and Pixel Shuffle (PS) layer in Model 4. Due to space restrictions and a smaller impact on the performance, we evaluate them together. For TULIP, we also apply Monte Carlo dropout [18] as a post-processing step, which helps to further reduce ghost points in between objects, which is reflected by the enhancement in IoU. For TULIP-L, we increase the model capacity, using four Swin-Transformer layers instead of three in both encoder and decoder, which demonstrates further performance boost on all metrics.
Inference Time: We evaluate the inference time on the KITTI dataset [19]. A single forward pass for TULIP-L takes 1.33s. For the Monte-Carlo Dropout we run the separate forward passes in a batch and calculate the filtered result, which increases the effective inference time of TULIP-L to 1.61s. This is slower than ILN [24] (1.14s) and LIDAR-SR [38] (1.16s) and comparable to Swin-IR [26] (1.63s).
@@ 4.3 Benchmark Results
Besides bilinear interpolation, which computes the interpolation weights from the four nearest neighbors for upsampling, we compete against the state-of-the-art LiDAR upsampling approaches Implicit LiDAR Network (ILN) [24] and LiDAR-SR [38]. We also compare against several image super-resolution works: SRNO  [46], and LIIF [10] learn implicit features for interpolation instead of 2D coordinates of new pixels while HAT [9] and Swin-IR [26] are pixel-based approaches that directly reconstruct high-resolution images from low-resolution feature embeddings. Similar to our work, they are also both based on Swin Transformer [29]. We train all methods on the specific datasets using their default parameters.
@@
ILN [24] 0.8592 0.5206 0.1855
TULIP (Ours) 0.7768 0.5291 0.1168
TULIP-L (Ours) 0.7748 0.5332 0.1032
Table 2: Quantitative comparison against state-of-the-art LiDAR and image super-resolution methods on simulation dataset. All methods are trained and evaluated on the same splits.
@@
TULIP (Ours) 0.4183 0.4134 0.1198
TULIP-L (Ours) 0.3678 0.4329 0.0971
@@
TULIP (Ours) 1.5382 0.3552 0.06560
TULIP-L (Ours) 1.5386 0.3620 0.06371
Table 3: Quantitative comparison against state-of-the-art LiDAR and image super-resolution methods on real-world dataset. All methods are trained and evaluated on the same splits.
@@
A
B
C
@@ 4.3.1 Simulation Results
Similar to prior works, we first evaluate the performance on noise-free, simulated data. We train and test all approaches on the dataset presented in  [24] that was recorded using the CARLA simulator and select the ground-truth range images of size $128\times 2048$ that capture a vertical FoV of ${}^{\circ}$. We use a (20699/2618) train/test split. We observe that the performance between image super-resolution approaches varies greatly, and LIIF performs the best among them. While LIIF [10] achieves a comparable MAE to the LiDAR upsampling methods, the 3D metrics are significantly worse, indicating that it is unsuitable for the underlying geometric input data. Nevertheless, LIIF [10] still outperforms the LiDAR upsampling method in LIDAR-SR. The interpolation-based approach in ILN shows decent results for the noise-free data, as indicated by their high IoU. Nevertheless, our approaches achieve the best performance in all evaluated metrics, with the strongest improvement in Chamfer Distance, which validates that our design is well suited for the geometric reconstruction in range images.
@@ 4.3.2 Real-World Results
KITTI: The KITTI dataset [19] was collected using a Velodyne HDL-64E LiDAR with a vertical FoV of ${}^{\circ}$ and a resolution of $64\times 1024$. For the image super-resolution approaches, we observe comparable results and trends on KITTI as on the simulated data in Sec.4.3.1. Even though LIDAR-SR [38] is designed for LIDAR upsampling and outperforms image super-resolution approaches in terms of MAE, it achieves the lowest IoU among all approaches, which can be explained by a large amount of invalid floating points between real objects. LIIF [10] achieves better CD and comparable IoU compared to ILN. We observe that the IoU of ILN [24] is drastically lower than on CARLA. The interpolation scheme in ILN is thus not suitable to properly handle the lower resolution data with noise from the real-world sensor. Our approaches significantly outperform all competing methods, especially in terms of IoU as it can better handle discontinuities between objects. As indicated in the visualizations in Fig.7, our approach predicts fewer invalid float points between objects, as well as fewer blurry edges, which results in much clearer point clouds.
DurLAR: The DurLAR dataset [25] was recorded using an Ouster OS1-128 LiDAR, that captures point clouds with a resolution of $128\times 2048$ with a ${}^{\circ}$ vertical FoV. Compared to CARLA and KITTI, we see a large performance drop for all approaches in MAE and IoU, which results from the larger maximum range and vertical FoV of the sensor as well as stronger sensor noise. We observe that all approaches designed for LiDAR upsampling clearly improve in Chamfer Distance compared to image super-resolution approaches. ILN and TULIP outperform all methods by a large margin in IoU. TULIP-L achieves the best 3D metrics, which validates that it can also handle the challenging DurLAR data better than prior works.
Range Analysis: As the density of LiDAR point clouds decreases with increasing distance from the sensor, we additionally compare the performance at different ranges for a more fine-grained evaluation. We observe that our approach achieves better or similar performance compared to all approaches in all metrics at all ranges. Up to 30m, TULIP clearly outperforms the baseline approaches. TULIP handles discontinuities between objects better than other approaches, which can be verified visually through sharper edges and fewer ”floating” points between objects. This is especially reflected in the high IoU. At ranges above 30m, our approach performs comparable to other methods. Above this range, LiDAR data is extremely sparse, meaning that points neighboring in the low-resolution range image are actually distant in the Euclidean space. It is, therefore, not feasible to infer meaningful information about the scene from such neighboring points at a high range, which is also represented in the poor performance of all approaches.

### v3

@@ 3.2.1 Tokenization
Tokenization is done by a patch partition layer that creates an initial feature embedding from the input range image. This initial feature representation significantly influences the network’s performance. Specifically, the selection of this layer determines how the inherent characteristics, patterns, and relevant information of the LiDAR range image data are encoded into a format that the network can effectively learn from. Most transformer-based RGB image super-resolution approaches utilize a relatively large patch size to tokenize the input [32, 10, 38]. These larger patches enable the construction of a global spatial context, facilitating an understanding of the overall structure and high-level features of the RGB image. These networks employ square-shaped patches and implement upsampling in both vertical and horizontal directions. In contrast, LiDAR upsampling primarily aims to enhance the vertical resolution of the input, given that range images possess properties that are vastly different from RGB images. Furthermore, attention across more pixels in range images is less useful than in RGB images, since there is almost no geometrical relation for the spatial contexts that are far from each other in the 2D range image space. Motivated by this, we propose two adjustments to effectively process range images within a Swin-Unet:
@@
Row-Based Patch Partition: Our model builds row patches with a dimension of $1\times 4$ for range image tokenization. The new patch geometry is designed to retain full vertical information while compressing horizontally. This aligns with our objective of conserving and extending vertical details in the range image. Besides that, row patches excel at capturing boundary discontinuities between distinct objects and the background within the scene. Moreover, the selection of the patch size strikes a balance between enhancing model performance and limiting model capacity. Smaller patches can capture more detailed local information, but they also increase the number of model parameters, which can slow down the model’s training process. Circular Padding (CP) in Horizontal Dimension: Instead of adding zero pixels padding, our model uses circular padding in the horizontal dimension. This avoids introducing artificial features and naturally matches the sensor model of a rotating 3D LiDAR. This preserves accurate neighbor relations along the edges, which would otherwise be compromised during the projection to the 2D range image. The Circular Padding (CP) also allows us to fit flexible input sizes without changing the patch geometry.
@@ 3.2.2 Non-Square window for local self-attention
Inside Swin Transformer, input feature maps are window-partitioned along width and height to compute the self-attention locally in each window. In addition to square-shaped RGB images, some recent work applies a square window for panoramic data as well [34, 61]. However, using square windows on range images is unfavorable. On one hand, small windows can focus more on attention within near areas but lead to increasing computational complexity and can struggle with the reconstruction of objects at a further distance in the scene. On the other hand, a larger square window can blend scene-related contexts at different scales, which can potentially degrade the network performance. In a prior study done by Eskandar et al. [19], they demonstrated a significant enhancement in range image super-resolution by separately upsampling the upper and lower portions of the data. In contrast to their approach of utilizing CNN-based shallower and deeper network branches for distinct feature extraction, our proposal involves the direct partitioning of the input using a rectangular window. This window geometry facilitates robust attention interaction among distant points through W-MSA and subtle cross-attention between objects at multiple scales via SW-MSA and hierarchical processing. In addition, range images, as typical panorama data, inherently contain abundant information along the horizontal direction. Therefore, directing attention more toward image width rather than height can aid the network in capturing a greater amount of geometrical information. At the same time, the computational complexity remains the same as using a large square window.
3.2.3 Further Adaptation and Refinement
Patch Unmerging (PU) and Pixel Shuffle (PS): Similar to the patch merging layer that is responsible for downsampling at the encoder stage, a mechanism is necessary within the decoder to upsample the feature maps. Conventional CNN-based approaches [53, 46] achieve upsampling through the transposed convolutional operation. As they tend to smooth out image sharpness [32], we reversed the operation of patch merging, re-arranging the channels of patches into their respective positions in a grid to upsample the spatial resolution of input feature maps, which we denote as patch unmerging. Furthermore, to reconstruct the final range image, we built the reconstruction head upon the upsampled feature maps. The implementation of the reconstruction module is based on the pixel shuffle layer [47]. Monte Carlo Dropout: We additionally use Monte Carlo Dropout [46] to filter unreliable points by thresholding the uncertainty. MC-Dropout executes several feed-forward passes during inference with different active dropouts, which yields a distribution of outputs. We refine the results by removing points with high output variance, as they often come down to noisy, invalid points.
@@ 4.1 Experimental Settings
Datasets: We conduct experiments on three different datasets, that include the two large-scale real-world datasets DurLAR [31] and KITTI [22] (Sec. 4.3.2), as well as on a dataset [28] that was generated using the CARLA simulator [17] (Sec. 4.3.1). We select test sequences that are recorded in different locations than the train set to avoid spatial overlap. We vertically downsample the high-resolution range images with a factor of four by skipping the respective lines to generate low-resolution input images. For ablation studies in Sec. 4.2, we select the KITTI dataset. Implementation Details: In all experiments, we use distributed processing with 4×4$\times$4 × GeForce RTX 2080 Ti. For training, we use a fixed batch size of eight per GPU for all datasets. We use the AdamW [37] optimizer with a weight decay factor of 0.01 and a base learning rate of 5e-4. Evaluation metrics: We evaluate the Mean Absolute Error (MAE) for all the pixels in the generated 2-D range images. Additionally, we assess the performance by considering the 3D points reconstructed by our neural networks. The Chamfer Distance (CD) measures the Euclidean distance between two point clouds. We also follow the approach in [28] to evaluate the volumetric occupancy similarity. To do so, the point clouds are voxelized using a voxel size of $0.1m$. A voxel is classified as occupied for each point cloud if it contains at least one point. We then calculate the Intersection-over-Unit (IoU) based on the occupancy.
@@ 4.2 Ablation Studies
Model 3 $1\times 4$ Rectangular $\times$ $\times$ 0.4248 0.4040 0.1218
Model 4 $1\times 4$ Rectangular $\checkmark$ $\times$ 0.4227 0.4084 0.1207
TULIP $1\times 4$ Rectangular $\checkmark$ $\checkmark$ 0.4185 0.4174 0.1207
TULIP-L $1\times 4$ Rectangular $\checkmark$ $\checkmark$ 0.3708 0.4329 0.0992
ViT-Unet* $1\times 4$checkmark $\checkmark$ 1.4484 0.2948 0.2665
Table 1: Ablation study results for 4×4$\times$4 × upsampling on KITTI. Input Resolution: $16\times 1024$, Output Resolution: $64\times 1024$. *We replaced Swin Transformer blocks with ViT blocks and followed the same training procedure as for TULIP.
@@
Window geometry: To verify the effectiveness of modifying window geometry for range images, we conduct experiments with square and rectangular windows (Tab. 1 Model 2-3. Based on the quantitative results, we can observe the following: reconstruction quality produced by local attention within large square windows is inferior to the rectangular window concerning both 2D and 3D evaluation metrics, while they result in the same computational complexity. The underlying reason for this performance difference lies in the perceptive field of the windows. A wider window is more likely to include discontinuities between objects at different scales. Accordingly, the network performs better at separating objects from each other, which can visually be validated by clearer surface boundaries in Fig. 4.
Further Adaptation and Refinement: Circular Padding (CP) helps maintain continuity and consistency at the edges of the panoramic images, but it has minor contributions to other areas. Patch Unmerging (PU), and Pixel Shuffle (PS) aim at upscaling the spatial resolution with rearrangement, avoiding information loss and parameter increase led by using additional de-convolutional layers. Although they are designed to improve efficiency rather than efficacy of the model, tested with those components (Model 4), a slight improvement in upsampling is still observable. Monte Carlo Dropout [20], as a post-processing step, helps to further reduce ghost points in between objects and we present more details in the Appendix. For TULIP-L, we increase the model capacity, using four Swin-Transformer layers instead of three in both encoder and decoder, which demonstrates an additional performance boost on all metrics.
Transformer Block: From the previous discussion, we infer that Swin Transformer [36] is more advantageous than ViT (Vision Transfromer) [18] in LiDAR upsampling. On one hand, prior works [21, 30, 35, 14] have pointed out the inferiority of ViT on smaller datasets due to lack of locality learning and non-overlapping attention. Relevant LiDAR datasets [22, 31] are generally small compared to RGB dataset [44]. On the other hand, by narrowing the patch size, the exponential increase of model capacity hinders sufficient training with ViT on range images, while Swin Transformer expands only linearly. To numerically show the distinction, we trained the network on the same dataset by replacing ViT blocks and applied the same patch size and additional components to the network for a fair comparison. In the last row of Tab. 1, it shows that TULIP significantly outperforms the one with ViT as the backbone.
@@ 4.3 Benchmark Results
Model MAE $\downarrow$ IoU $\uparrow$ CD $\downarrow$
@@
ILN [28] 0.8592 0.5006 0.1855
TULIP (Ours) 0.7699 0.5152 0.1028
TULIP-L (Ours) 0.7539 0.5301 0.1001
@@
TULIP (Ours) 0.4185 0.4174 0.1207
TULIP-L (Ours) 0.3708 0.4329 0.0992
@@
TULIP (Ours) 1.5432 0.3562 0.06484
TULIP-L (Ours) 1.5592 0.3654 0.06346
Table 2: Quantitative comparison against state-of-the-art LiDAR and image super-resolution methods on different datasets. All methods are trained and evaluated on the same splits.
Besides bilinear interpolation, which computes the interpolation weights from the four nearest neighbors, we evaluate against the state-of-the-art LiDAR upsampling approaches Implicit LiDAR Network (ILN) [28] and LiDAR-SR [46]. We also compare against several image super-resolution works: SRNO  [54], and LIIF [11] learn implicit features for interpolation instead of 2D coordinates of new pixels while HAT [10] and Swin-IR [32] are pixel-based approaches that directly reconstruct high-resolution images from low-resolution feature embeddings. Similar to our work, they are based on Swin Transformer [36]. We train all methods on the specific datasets using their default parameters.
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
A B C
@@ 4.3.1 Simulation Results
Similar to prior works, we first evaluate the performance on noise-free, simulated data. We train and test all approaches on the dataset presented in  [28] that was recorded using the CARLA simulator and select the ground-truth range images of size $128\times 2048$ that capture a vertical FoV of 30. We use a (20699/2618) train/test split. We observe that the performance between image super-resolution approaches varies greatly, and LIIF performs the best among them. While LIIF [11] achieves a comparable MAE to the LiDAR upsampling methods, the 3D metrics are significantly worse, indicating that it is unsuitable for the underlying geometric input data. Nevertheless, LIIF [11] still outperforms the LiDAR upsampling method in LIDAR-SR. The interpolation-based approach in ILN shows decent results for the noise-free data, as indicated by their high IoU. Nevertheless, our approaches achieve the best performance in all evaluated metrics, with the strongest improvement in Chamfer Distance, which validates that our design is well suited for the geometric reconstruction in range images.
@@ 4.3.2 Real-World Results
KITTI: The KITTI dataset [22] was collected using a Velodyne HDL-64E LiDAR with a vertical FoV of 26.8 and a resolution of $64\times 1024$. For the image super-resolution approaches, we observe comparable results and trends on KITTI as on the simulated data in Sec.4.3.1. Even though LIDAR-SR [46] is designed for LIDAR upsampling and outperforms image super-resolution approaches in terms of MAE, it achieves the lowest IoU among all approaches, which can be explained by a large amount of invalid floating points between real objects. LIIF [11] achieves better CD and comparable IoU compared to ILN. We observe that the IoU of ILN [28] is drastically lower than on CARLA. The interpolation scheme in ILN is thus not suitable to properly handle the lower resolution data with noise from the real-world sensor. Our approaches significantly outperform all competing methods, especially in terms of IoU as it can better handle discontinuities between objects. As indicated in the visualizations in Fig.5, our approach predicts fewer invalid float points between objects, as well as fewer blurry edges, which results in much clearer point clouds.
DurLAR: The DurLAR dataset [31] was recorded using an Ouster OS1-128 LiDAR, that captures point clouds with a resolution of $128\times 2048$ with a 45 vertical FoV. Compared to CARLA and KITTI, we see a large performance drop for all approaches in MAE and IoU, which results from the larger maximum range and vertical FoV of the sensor as well as stronger sensor noise. We observe that all approaches designed for LiDAR upsampling clearly improve in Chamfer Distance compared to image super-resolution approaches. ILN and TULIP outperform all methods by a large margin in IoU. TULIP-L achieves the best 3D metrics, which validates that it can also handle the challenging DurLAR data better than prior works.
Range Analysis: As the density of LiDAR point clouds decreases with increasing distance from the sensor, we additionally compare the performance at different ranges for a more fine-grained evaluation. We observe that our approach achieves better or similar performance compared to all approaches in all metrics at all ranges. Up to 30m, TULIP clearly outperforms the baseline approaches. TULIP handles discontinuities between objects better than other approaches, which can be verified visually through sharper edges and fewer "floating" points between objects. This is especially reflected in the high IoU. At ranges above 30m, our approach performs comparable to other methods. Above this range, LiDAR data is extremely sparse, meaning that points neighboring in the low-resolution range image are actually distant in the Euclidean space. It is, therefore, not feasible to infer meaningful information about the scene from such neighboring points at a high range, which is also represented in the poor performance of all approaches.